# WeQA: A Benchmark for Retrieval Augmented Generation in Wind Energy Domain

### v2

@@ 4 Methodology
Generate {num} questions given the content provided in the following paragraph. Restrict the type of questions to {question type} questions.
{Text chunk to generate the questions.}
You can generate similar questions (but not limited) to sample questions provided below.
{List of sample questions}
Hybrid text chunks.  A problem with the aforementioned approach is that a significant number of questions are generated on a single sentence basis. This is obtained by substituting the subject or object of a sentence with a ‘wh’ word. These generated questions are meaningful when we consider question types such as ‘closed’, ‘open’, or ‘recall’, where the answers can be a single sentence from the provided text chunk. However, ‘process’, ‘evaluation’, and ‘comparison’ type questions of enhanced quality require the answer to be inferred from a larger portion of the given text chunk. To this end, first we use GPT-4 to summarize the entire text chunk (consisting of more than 15 sentences) into a summarized text chunk (consisting 5-8 sentences) using a prompt as follows:
@@
You are a smart assistant. Can you summarize this input paragraph within {num} bullet points. Return the summarized text.
Input paragraph: “‘ {Text chunk to summarize} ”’
Thereafter, we use GPT-4 with appropriate prompts to generate questions from these summarized text chunks using the previous hybrid prompt along with the list of sample questions. Here, we show an example question generated using this approach. We include the summary text chunk generated by GPT-4 in Figure 3 and highlight the text in red color, from which the answer for the ‘comparison’ type question is retrieved.
Question: How does the proportion of cultivated cropland within the Pleasant Ridge Wind Resource Area (PRWRA) compared to the proportion of developed areas?
Answer: Cultivated cropland covers $92.3\%$ of the PRWRA while developed areas cover $5.1\%$.
@@
GPT-4 as Evaluator Gemini 1.5 Pro as Evaluator
@@
Table 2: Performance of the models on the WeQA benchmark scored using the RAGAS framework across evaluators. The "Prec." and "Rec." mean Context Precision and Context Recall respectively, while "Type" refers to the Question Type. The best performance for each question type per evaluator is highlighted in bold.
Questions from tables.  Another important aspect of benchmarking RAG models in the domain of research articles and reports is to evaluate their performance in retrieving information from tables. Tables are important contents inside research documents and often contain useful summaries of the entire documents.
Generate {num} questions given the table provided in HTML format in the following paragraph? Generate the questions keeping in mind that the caption of the table is “‘ {Table caption obtained from document.} ”’ Restrict the questions such that the answers are only from the provided table in the html format. For each question, return 3 lines: question/ answer/ proof. Make sure there are no newline characters in the proof.
Input table:
“‘ {Table in HTML format extracted from document} ”’
Table 1 shows a table from the report Invenergy (2014) and we generate questions from this table as follows.
Question: What is the acreage of Cultivated Crops within the Pleasant Ridge Project Area based on the National Land Cover Database in May of 2014?
Answer: The acreage of Cultivated Crops within the Pleasant Ridge Project Area is 55,946 acres.
Proof: The table entry under the “Habitat” column for “Cultivated Crops” corresponds with the entry under the “Acres [Hectares]” column that reads “55,946[22,641]”
5 Results and Discussion
We evaluate three RAG-based LLM s, namely GPT-4, Gemini, and Claude, on our WeQA benchmark. The $\mathsf{RAGAS}$ framework is employed for this evaluation, utilizing an evaluator LLM to assess the models’ performance. The assessment includes metrics such as answer correctness, context precision, and context recall, providing a comprehensive understanding of each model’s capabilities in retrieving and generating accurate information from the given context. In our case, we have used GPT-4 and Gemini-1.5Pro as choices for the evaluator LLM s. Figure 4 presents the answer correctness score, while context precision and context recall depicted in Table 2 show the ability of the models to retrieve the context accurately.
Observation 1
Observation 2
There is an alignment in evaluations made by the two evaluator LLM s used within the $\mathsf{RAGAS}$ framework, particularly visible for ‘closed’ type questions.
This similarity arises because the answers to these questions are objective (‘yes’ or ‘no’), leading to equivalent correctness evaluations by both models. Although there are some mismatches in the evaluations made by the two evaluator models, the number of these discrepancies is insignificant compared to the number of matching evaluations.
Figure 4: Answer correctness scores computed using the RAGAS scoring framework with GPT-4 and Gemini-1.5Pro as evaluator models for response generated by all three models used.
Figure 5 displays the confusion matrix illustrating the evaluations made by the two evaluator LLM s (GPT-4 and Gemini-1.5Pro) on the responses provided by the RAG-based Claude and GPT-4 models to the benchmarking questions. In this context, a true positive occurs when the LLM evaluator correctly identifies the model response as matching the ground truth. Conversely, a false positive arises when the LLM evaluator incorrectly states that the model response matches the ground truth, while it does not. This matrix helps visualize the accuracy and reliability of the evaluations conducted by the LLM s, when used within the $\mathsf{RAGAS}$ framework. We note that majority of evaluations made by either LLM evaluator matches the actual evaluation which indicates that both of them are reliable.
Observation 3
From this observation, we conclude that RAG-based models generate more accurate subjective responses to ‘open’ questions than objective (‘yes’ or ‘no’) responses for ‘closed’ questions. This suggests that these models perform better when tasked with generating detailed, context-rich answers rather than simple, binary ones, highlighting their strength in handling nuanced and complex queries.
Figure 5: Confusion Matrix for evaluations by LLM evaluators on responses from Claude (top) and GPT-4 (bottom) models
Observation 4
Figure 6: Answer correctness (top left), answer similarity (top right), context precision (bottom left) and recall (bottom right) scores across different evaluator and generator models.
Observation 5
Observation 6
Figure 6 shows the scores computed by the evaluators for the responses generated by the three RAG-based LLM s. The Gemini-1.5Pro evaluator tends to rate high scores even when the LLM s refuse to answer. An example is listed below:
Question: In the HTML table that estimates the annual number of bird collisions at different percentages of avoidance, what is the estimated number of collisions per year at $98.0\%$ avoidance?
Expected answer: The estimated number of collisions per year at $98.0\%$ avoidance is $152$.
Gemini generated answer: This question cannot be answered from the given source. While the text discusses bird collisions with wind turbines, it does not provide any data on estimated collision numbers at different avoidance rates.
GPT-4 score: 0.194
Gemini-1.5Pro score 0.813
Conversely, in instances where the LLMs generate correct answers, Gemini-1.5Pro has been observed to evaluate them as incorrect. An example is detailed below:
Question: Who is the GIS Technician in the ‘STUDY PARTICIPANTS’ table?
Expected answer: JR Boehrs
Gemini generated answer: Saif Nomani JR Boehrs was the GIS Technician.
GPT-4 score: 0.703
Gemini-1.5Pro score: 0.200
6 Conclusion
In conclusion, this paper presents a versatile framework for evaluating the performance of RAG-based LLM s across various question types and document sections. We showcase this by introducing a hybrid, automated question-generation method that ensures comprehensive coverage of both objective and subjective queries, and implement this for the use case of wind energy related document and present the WeQA benchmark, which is a first of its kind benchmark in wind energy domain. However, the usefulness of our work goes beyond this niche domain as our approach is domain-agnostic, meaning it can be used for creating benchmark for any domain. Additionally, our use of the RAGAS scoring framework comes with multiple benefits; it allows for a thorough evaluation of model performance, offering a holistic assessment of LLM capabilities, while also having the advantage of being easy for other researchers to adapt this approach for their own work.
7 Limitations
A limitation of the proposed framework is that the automatic method of generating questions often produces queries that are too specific to the document from which they were derived. When these questions are posed to an LLM with a large document corpus, the model may struggle to respond accurately, necessitating the filtering of ambiguous questions to ensure relevance and clarity. Additionally, the $\mathsf{RAGAS}$ scoring framework, which relies on LLM s as evaluators, introduces uncertainty in performance metrics, as different LLM evaluators may score responses differently. While comparisons can be made for questions with objective responses, evaluating and comparing subjective responses across different LLM s remains challenging and less consistent.
8 Ethical Considerations
9 Acknowledgment
This research was sponsored by US DOE Wind Energy Technologies Office (WETO) and was done at Pacific Northwest National Laboratory, a multi-program national laboratory operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RLO1830. This article has been cleared by PNNL for public release as PNNL-SA-203893.

### v3

@@ Appendix A Prompts used to generate QA pairs using Hybrid Prompt Approach
Prompt with placeholder Generate {number of questions} questions given the content provided in the following paragraph. Restrict the type of questions to {question type} questions. {Text chunk from document section}
We curate the generated questions, where domain experts manually identify the questions which are best suited for the purpose of benchmarking LLM s. We perform this process for each type of question, so that we include particular grammatical structures for each question type. Thereafter, we use these curated high-quality questions as few-shot examples to regenerate questions using the automatic question generation framework. The updated prompt along with the placeholders looks as follows:
Prompt with placeholder Generate {number of questions} questions given the content provided in the following paragraph. Restrict the type of questions to {question type} questions. {Text chunk from document section} You can generate similar questions (but not limited) to sample questions provided below. {Sample question 1} {Sample question 2} {Sample question 3}
Appendix B Prompts used to generate QA pairs using Hybrid Context Approach
We use the following prompt to summarize a document section from which the questions are to be generated.
@@ Appendix B Prompts used to generate QA pairs using Hybrid Context Approach
Prompt with placeholder You are a smart assistant. Can you summarize this input paragraph within {number of points} bullet points. Return the summarized text. Input paragraph: {Text chunk from document to summarize}
Thereafter, we use the earlier prompt to generate questions from this summarized text chunk. We add the few-shot example questions which are identified by the domain experts for each question type.
Prompt with placeholder Generate {number of questions} questions given the content provided in the following paragraph. Restrict the type of questions to {question type} questions. {Summarized text chunk from document section} You can generate similar questions (but not limited) to sample questions provided below. {Sample question 1} {Sample question 2} {Sample question 3}
Appendix C Prompts used to generate QA pairs from tables
We extract the tabular data from documents as HTML objects in the filtered JSON schema. We use the following prompt to generate question-answer pairs from the tabular data.
Prompt with placeholder Generate {number of questions} questions given the table provided in HTML format in the following paragraph? Generate the questions keeping in mind that the caption of the table is {Table caption obtained from document.} Restrict the questions such that the answers can be retrieved from the provided table in the HTML format. For each question, return 3 lines: question/ answer/ proof. Make sure there are no newline characters in the proof. Input table:{Table in HTML format extracted from document}
We show an example QA pair generated from a table obtained from a document Invenergy (2014). Table 2 shows the table from the document for reference. An example QA-pair generated from this table is provided here.
LLM generated question-answer pair Question: What is the acreage of Cultivated Crops within the Pleasant Ridge Project Area based on the National Land Cover Database in May of 2014? Answer: The acreage of Cultivated Crops within the Pleasant Ridge Project Area is 55,946 acres. Proof: The table entry under the “Habitat” column for “Cultivated Crops” corresponds with the entry under the “Acres [Hectares]” column that reads “55,946[22,641]
@@ Appendix C Prompts used to generate QA pairs from tables
Appendix D Context Recall and Context Precision
We utilize $\mathsf{RAGAS}$ context recall and precision metrics to evaluate the retrieval performance of our RAG-based systems, where context recall measures the proportion of relevant information successfully retrieved from the knowledge base, and context precision assesses the relevance of the retrieved context to the given query. In our setup, we employ semantic similarity-based retrieval using vector embeddings, where ‘relevant context’ is defined as text chunks or the document sections that contain information necessary to answer the posed questions.
GPT-4 as Judge Gemini 1.5 Pro as Judge
@@ Appendix D Context Recall and Context Precision
Table 3: Performance of the models on the WeQA benchmark scored using the RAGAS framework across judge LLM s. The "Prec." and "Rec." mean Context Precision and Context Recall respectively, while "Type" refers to the Question Type. The best performance for each question type per judge LLM is highlighted in bold.
Appendix E Judge LLM Evaluation Analysis Through Confusion Matrices
To assess the reliability and accuracy of LLM s as judges within the $\mathsf{RAGAS}$ evaluation framework, we conduct a detailed analysis using confusion matrices for closed-type questions where binary (‘yes’/‘no’) responses can be objectively compared against ground truth answers. This analysis is particularly crucial for validating the trustworthiness of automated evaluation systems in benchmarking scenarios.
Methodology for evaluation. We evaluate two judge LLM s—GPT-4 and Gemini-1.5Pro—by comparing their assessments of RAG-based model responses (Claude and GPT-4) against manually verified ground truth labels for closed-type questions. The confusion matrix framework allows us to quantify four key evaluation scenarios:
  • True Positive (TP): The judge LLM correctly identifies that the model response matches the ground truth answer.
  • False Positive (FP): The judge LLM incorrectly states that the model response matches the ground truth when it does not
  • True Negative (TN): The judge LLM correctly identifies that the model response does not match the ground truth answer
  • False Negative (FN): The judge LLM incorrectly states that the model response does not match the ground truth when it actually does
Analysis of judge LLM performance. The confusion matrices reveal that the majority of evaluations made by both judge LLMs align with the actual ground truth evaluations, demonstrating their reliability as automated evaluators. Specifically, both GPT-4 and Gemini-1.5Pro exhibit high accuracy rates in distinguishing correct from incorrect responses, with minimal discrepancies in their assessment capabilities.
Cross-judge agreement. Additionally, we observe substantial agreement between the two judge LLM s, suggesting consistency in evaluation standards across different model architectures. This cross-validation approach enhances the robustness of our evaluation methodology and provides confidence in the reliability of automated assessment within specialized domain benchmarks like WeQA.